# Evaluating the Capability of ChatGPT on Ancient Chinese

### 23 Dec 2023 v1

@@ 4.1 Assessing the Efficacy of ChatGPT in the Translation of Ancient Chinese Texts
Table 3: The BERT-Scores, precision, recall and F1 score, of ancient-to-modern Chinese translation from ChatGPT under four settings: one sentence (1-sent.), three sentences (3-sent.), five sentences (5-sent.), and eight sentences (8-sent.) per query.
@@
Table 3 shares the same structure as Table 2, but displays the BERT-Scores, precision, recall, and F1 scores associated with the translation of ancient to modern Chinese by ChatGPT. From this table, ChatGPT performs the best when translating three consecutive ancient Chinese sentences (3-sent.), achieving 78.5% F1 score, followed by 5-sent. and 8-sent. This finding is consistent to BLUE scores in Table 2.
Table 4: The groundtruth (GT) and translated texts from ChatGPT under four settings: one sentence (1-sent.), three sentences (3-sent.), five sentences (5-sent.), and eight sentences (8-sent.) per query.
@@
GT [CN]:孙秀既恨石崇不把绿珠送给他,又不满潘岳从前对自己不礼貌。
[EN]:Sun Xiu both resented Shi Chong for not giving her the Green Pearl, and
was dissatisfied with Pan Yue for his past discourteous behavior towards her.
@@
[EN]:Sun Xiu both held resentment towards Shi Chong for not marrying Green
Pearl, and also lamented over Pan Yue’s past discourteous behavior towards him.
@@
[EN]:Sun Xiu felt resentment towards Shi Chong for not giving the Green Pearl
@@
[EN]:Sun Xiu both resented Shi Chong for his inability to be with Green Pearl and
@@
[EN]:Sun Xiu harbored resentment towards Shi Chong for not marrying Green Pearl,
@@
To gain deep insights, we select and analyze a specific sentence as a case study. The original ancient Chinese sentence is ”孙秀既恨石崇不与绿珠,又憾潘岳昔遇之不以礼。”, which means that Sun Xiu both resented Shi Chong for not giving her the Green Pearl and was also upset that Pan Yue did not treat her with respect in the past. The groundtruth modern Chinese and translated texts from four settings are exhibited in Table 4. To aid readers understand the translated Chinese, we also provide an English version for all methods. As shown in Table 4, The translated modern Chinese texts from ChatGPT missed the Green Perl part except feeding three sentences (3-sent.), as highlighted by the underlines.
@@
4.2 Assessing the Efficacy of People Name Recognition in Ancient Chinese using ChatGPT
@@
People name recognition task can assess the capability of models on understanding ancient Chinese. For the people name recognition task, a total of 300 sentences are randomly selected, and a few examples are shown in Table 1.
@@
Table 5: Comparative performance of ChatGPT (GPT-3.5-Turbo) and Jieba in recognizing people names from 300 sentences in Shi Shuo Xin Yu.
ChatGPT Jieba
Precision 87.01% 44.63%
Recall 70.16% 52.94%
F1 score 77.58% 48.43%
@@
Table 5 presents the performance of ChatGPT (GPT-3.5-Turbo) and Jieba in recognizing people names from ancient Chinese texts. This evaluation employed 300 sentences sourced from Shi Shuo Xin Yu. From the table, ChatGPT performs reasonably well on the personal name recognition task, achieving 77.58% F1 score, much better than Jieba.
@@
ChatGPT also faces many challenges in recognizing personal names from ancient Chinese texts, because understanding ancient Chinese sentences rely heavily on the context of the sentence. Sometimes a sentence may omit the name of the person, or replace it with other names.
5 Conclusion and Discussion
Through experiments on the capacity of ChatGPT on ancient-to-modern Chinese translation and personal name recognition, we find that there are still much space for improvement. One possible explanation might be that ChatGPT is mainly pre-trained on English corpora with limited amount of Chinese corpora.
@@ 5 Conclusion and Discussion
While the proficiency of ChatGPT in ancient Chinese remains somewhat restricted, we assert that engagements of deeper research and enhancements - coupled with an abundant corpus of ancient Chinese and specialized knowledge - can ameliorate its proficiency in comprehending and translating ancient Chinese text. This improvement can potentially expand the possibilities for studying ancient history and culture, along with fostering the preservation and progression of traditional cultural heritage.

### 23 Feb 2024 v2

@@ IV-A Assessing the Efficacy of ChatGPT in the Translation of Ancient Chinese Texts
TABLE III: The BERT-Scores, precision, recall and F1 score, of ancient-to-modern Chinese translation from ChatGPT under four settings: one sentence (1-sent.), three sentences (3-sent.), five sentences (5-sent.), and eight sentences (8-sent.) per query.
@@
Figure 3: The BERT-scores of ancient-to-modern Chinese translation when ChatGPT is fed with 1, 3, 5, and 8 consecutive sentences.
Table III shares the same structure as Table II, but displays the BERT-Scores, precision, recall, and F1 scores associated with the translation of ancient to modern Chinese by ChatGPT. From this table, ChatGPT performs the best when translating three consecutive ancient Chinese sentences (3-sent.), achieving 78.5% F1 score, followed by 5-sent. and 8-sent. This finding is consistent to BLUE scores in Table II.
To visualize the effects of input sentence number on the translation quality, Fig. 3 presents the line plots of BERT scores versus input sentence number in the prompt. From this figure, it can be observed that BERT scores (recall, precision and F1 score) peak at 3. This can be explained by the fact that ancient Chinese is complex and context is very important for understanding the pronouns and references. But long context may further confuse the model, so the performance is best at a moderate length of 3 sentences.
TABLE IV: The groundtruth (GT) and translated texts from ChatGPT under four settings: one sentence (1-sent.), three sentences (3-sent.), five sentences (5-sent.), and eight sentences (8-sent.) per query.
@@
GT [CN]:孙秀既怨恨石崇不肯送出绿珠,又不满潘岳从前对自已不礼貌。
[EN]:Sun Xiu both resented Shi Chong for not giving him Zhu Lv (name of a woman),
and was dissatisfied with Pan Yue for his past discourteous behavior towards him.
@@
[EN]:Sun Xiu both held resentment towards Shi Chong for not marrying Zhu
Lv, and also lamented over Pan Yue’s past discourteous behavior towards him.
@@
[EN]:Sun Xiu felt resentment towards Shi Chong for not giving the Zhu Lv
@@
[EN]:Sun Xiu both resented Shi Chong for his inability to be with Zhu Lv and
@@
[EN]:Sun Xiu harbored resentment towards Shi Chong for not marrying Zhu Lv,
@@
To gain deep insights, we select and analyze a specific sentence as a case study. The original ancient Chinese sentence is “孙秀既恨石崇不与绿珠,又憾潘岳昔遇之不以礼。”, which means that Sun Xiu both resented Shi Chong for not giving him Zhu Lv (name of a woman), and was dissatisfied with Pan Yue for his past discourteous behavior towards him. The groundtruth modern Chinese and translated texts from four settings are exhibited in Table IV. To aid readers understand the translated Chinese, we also provide an English version for all methods. As shown in Table IV, The translated modern Chinese texts from ChatGPT is closest to the groundtruth when feeding three sentences (3-sent.). The wrong points are highlighted by the underlines.
@@
IV-B Assessing the Efficacy of People Name Recognition in Ancient Chinese using ChatGPT
@@
People name recognition task can assess the capability of models on understanding ancient Chinese. For the people name recognition task, a total of 300 sentences are randomly selected, and a few examples are shown in Table I.
@@
TABLE V: Comparative performance of ChatGPT (GPT-3.5-Turbo) and Jieba in recognizing people names from 300 sentences in Shi Shuo Xin Yu.
Precision Recall F1 score
Zero-shot performance
Jieba 44.63% 52.94% 48.43%
ChatGPT 72.66% 55.11% 62.68%
ChatGPT Few-shot performance
One-shot 72.76% 55.30% 62.90%
Five-shot 72.86% 55.68% 63.12%
Ten-shot 73.80% 56.82% 64.21%
Figure 4: The performance of ChatGPT on person name recognition versus the number of examples in the prompt.
@@
Table V presents the performance of ChatGPT (GPT-3.5-Turbo) and Jieba in recognizing people names from ancient Chinese texts in few-shot settings. This evaluation employed 300 sentences sourced from Shi Shuo Xin Yu. From this table, on the ancient person name recognition task, ChatGPT significantly outperforms Jieba in zero-shot setting, achieving 62.68% F1 score. But it still has much space to improve.
For the few-shot setting, the performance of ChatGPT gradually improves as the number of demonstrations increases in the prompt. Fig. 4 exhibits this increasing trend as well. As the number of instances increases from 0 to 10, the F1 score rises from 0.628 to 0.642.
IV-C Limitations
In our experiments, we did not systematically design and test prompt templates. We just pick the one that performs well on a few examples. In the future, we will investigate how to design efficient prompt templates for in-context learning.
@@ IV-C Limitations
During our experiments, we find that ChatGPT faces many challenges in recognizing personal names from ancient Chinese texts. Fully understanding ancient Chinese sentences rely heavily on the context of the sentence. Sometimes a sentence may omit the name of the person, or replace it with other names. Therefore, feeding multiple sentences in the prompt of ChatGPT. may enhance the performance as well. We leave it for future research.
V Conclusion and Discussion
@@
Through experiments on the capacity of ChatGPT on ancient-to-modern Chinese translation and person name recognition, we find that there are still much space for improvement. One possible explanation might be that ChatGPT is mainly pre-trained on English corpora with limited amount of ancient Chinese corpora. While the proficiency of ChatGPT in ancient Chinese remains somewhat restricted, we assert that engagements of deeper research and enhancements - coupled with an abundant corpus of ancient Chinese and specialized knowledge - can ameliorate its proficiency in comprehending and translating and understanding ancient Chinese text. This improvement can potentially expand the possibilities for studying ancient history and culture, along with fostering the preservation and progression of traditional cultural heritage.